OpenStudy (anonymous):

Solve the absolute value equation. Graph the solution. 2|3x – 5| – 8 = 8 (1 point)

OpenStudy (anonymous):

simplify it down first: 2|3x – 5| – 8 = 8 2|3x-5| = 16 |3x - 5| = 32 would you know how to proceed from here? what is the meaning of an 'absolute value' equation?

OpenStudy (anonymous):

I have no idea how to get to that point. I am really struggling with this.